OpenStudy (anonymous):

The legs of a right triangle are lengths x and x√3. The cosine of the smallest angle of the triangle is _____. 1/2 √3 (√3)/2 2√3

OpenStudy (anonymous):

help please

OpenStudy (noorik.):

do you know the law of cosine?

OpenStudy (anonymous):

yes cos= adj/hypo

OpenStudy (anonymous):

@NooriK.

OpenStudy (noorik.):

the law is \[\cos a = \frac{ B^2 + C^2 - A^2 }{ 2bc }\] A,B and C are the legs of the triangle a,b and c are the angles of the triangle

OpenStudy (anonymous):

okay so how can I use the info they gave to solve this equation?

OpenStudy (noorik.):

sorry i got drifted away, it's like you have said cos = adj / hyp then you would divide \[\cos = \frac{ x }{ x \sqrt{3} }\]

OpenStudy (anonymous):

you cant have root 3 on the bottom so its x root 3/3x?

OpenStudy (anonymous):

or do the x's cancel out?

OpenStudy (noorik.):

in the question you haven't told which one is the adj and which is the hyp, so i suppose it would be the opposite \[\cos = x \sqrt{3}/x\] but in other problems you have to draw the triangle to solve the problem, or you'll get wasted, the i'm right now XD

OpenStudy (anonymous):

wait what?

OpenStudy (noorik.):

you have to draw the triangle like then it would be solved or you'll get lost

OpenStudy (noorik.):

without drawing it would be hard to solve

OpenStudy (anonymous):

can you solve it?

OpenStudy (noorik.):

\[\cos = adj/hypo = x \sqrt{3}/x = \sqrt{3}\]

OpenStudy (anonymous):

so the x's cancel out giving us root 3 as the answer?

OpenStudy (noorik.):

indeed